OpenStudy (coledrf):

what is the simplified form of the following expression 1 over 4 (16-12x)

OpenStudy (shamil98):

\[\frac{ 16-12x }{ 4 } = \frac{ 4(4 - 3x) }{ 4 } = \frac{ \cancel 4 (4-3x) }{ \cancel 4 } = 4-3x\]

OpenStudy (coledrf):

wow how do i do that

OpenStudy (shamil98):

1/4(16-12x) you take the greatest common factor (GCF)

OpenStudy (coledrf):

what is the sum 7/4 + (-1/5)

OpenStudy (shamil98):

\[\frac{ 7 }{ 4 } - \frac{ 1 }{ 5 } = \frac{ 35 }{ 20 } - \frac{ 4 }{ 20 }\] just find a common denominator and subtract

OpenStudy (shamil98):

\[\frac{ 35 }{ 20 } - \frac{ 4 }{ 20 } = \frac{ 35-4 }{ 20 }\]